OpenStudy (kellygirl34):

Given the following functions f(x) and g(x), solve (f + g)(5) and select the correct answer below: (1 point) f(x) = 8x − 13 g(x) = x − 9

OpenStudy (arthur326):

The notation \((f+g)(5)\) is the same as \(f(5) + g(5)\). What is \(f(5)\) and \(g(5)\)?

OpenStudy (calculusxy):

(f + g)(5) also means f(5) + g(5) so find what f(5) = 8(5) - 13 and what g(5) = 5 - 9 then add them together

OpenStudy (calculusxy):

@kellygirl34 Did you find the answer?

OpenStudy (kellygirl34):

the options are 23, 27, 31, and 40

OpenStudy (kellygirl34):

i found the answer. It is 23

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\)\(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)